# Physical meaning of non-extensive term in Massieu functions

### v1

@@ I Introduction
We present a general approach to demonstrate that those differences can be written in terms of thermodynamic fluctuations, that, in turn, can be obtained from thermodynamic quantities. This means that the thermodynamic equation of state contains the information needed to calculate non-extensive terms of statistical mechanics quantities. Or, the other way around, non-extensive terms of statistical mechanics quantities have information about thermodynamic fluctuations. Our analysis is restricted to systems large enough so finite size effects, analyzed in, for example, Hill (1964); Bedeaux et al. (2020), other than the logarithmic term, can be neglected.
The difference $O(\ln N)$ is present in the non-ideal and ideal (negligible interactions) cases. Excess extensive quantities are obtained when the ideal value is subtracted. It can be seen that thermodynamic and statistical mechanics excess quantities differ by a term that is $O(N^{0})$. For example, the excess free energy of a system of particles, calculated using the partition function, includes a non-extensive term $O(N^{0})$; and the excess chemical potential has a non-extensive term $O(N^{-1})$. This term was calculated using the Widom insertion formula in Ref. Siepmann et al. (1992), where the result is presented as a finite size correction to numerical calculation of the excess chemical potential in a system with periodic boundary conditions. An equivalent expression was derived with a different procedure in Ref. Di Muro and Hoyuelos (2021), where it was shown that the non-extensive term is relevant for the evaluation of transition rates and, in turn, of the diffusion coefficient. Actually, the relationship between self-diffusion and collective diffusion coefficients, known as one of the the Darken equations Darken (1948), was derived in Di Muro and Hoyuelos (2021) using the non-extensive term of the excess chemical potential. Here we present a general approach, valid for any Massieu function, that includes the mentioned results; that is, it is not restricted to the analysis of the free energy.
@@ II Basic equations and notation
where $\Omega(\vec{X})$ is the number of microstates that are consistent with the macrostate $\vec{X}$. Units such that $k_{B}=1$ are used in order to simplify the notation; in this context, temperature has energy units. We use a tilde ($\tilde{S}$) in order to identify statistical mechanics quantities. The thermodynamic entropy, homogeneous of order 1 in $N$, is $S(\vec{X})$. The difference is $\tilde{S}-S=O(\ln N)$.
First, we present the definition of Massieu functions in their thermodynamic (TH) and statistical mechanics (SM) versions, $\Psi$ and $\tilde{\Psi}$, for a specific situation: a particle sub-system of volume $V$ that is part of a larger complex; the complement of the sub-system plays the role of a reservoir that imposes a temperature $T$ and a chemical potential $\mu$. Local fluctuations in $U$ and $N$ are present, and $V$ is a constant parameter by definition. We consider that the vector $\vec{X}=(U,N)$ contains the extensive variables that fluctuate, but there can be other variables, like $V$ in this case, that are fixed by external constraints. TH Massieu functions are obtained from successive applications of Legendre transforms, starting from the entropy. For example:
$\Psi(U,N)$ $=S(U,N)$
$\Psi(\tfrac{1}{T},N)$ $=S(U,N)-\tfrac{1}{T}U$ (2)
$\Psi(\tfrac{1}{T},\tfrac{\mu}{T})$ $=S(U,N)-\tfrac{1}{T}U+\tfrac{\mu}{T}N.$
Different Massieu functions are identified by their variables. We have that $\Psi(\tfrac{1}{T},N)=-F/T$, where $F$ is the Helmholtz free energy, and $\Psi(\tfrac{1}{T},\tfrac{\mu}{T})=PV/T$, where $P$ is the pressure. A fundamental hypothesis of thermodynamics is that thermodynamic potentials, and TH Massieu functions, are homogeneous functions of order 1 in $N$; that is, they are extensive quantities.
Now we turn to a statistical mechanics description. We have to specify how to calculate Massieu functions, $\tilde{\Psi}$, in this context. It is shown below that, instead of the Legendre transform, the mathematical transformation that generates $\tilde{\Psi}$ starting from $\tilde{S}$ is the Laplace transform. Quantities with tilde, in the statistical mechanics context, contain non-extensive terms, and we are interested in the difference $\tilde{\Psi}-\Psi$ at order $\ln N$.
The TH Massieu functions $\Psi(\tfrac{1}{T},N)$ and $\Psi(\tfrac{1}{T},\tfrac{\mu}{T})$, equal to $-F/T$ and $PV/T$, correspond to the logarithm of the canonical and grand canonical partition functions, $\mathcal{Z}(\tfrac{1}{T},N)$ and $\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})$, respectively. The equality holds when non-extensive terms are neglected in the thermodynamic limit. The SM Massieu functions are given by
$\tilde{\Psi}(\hat{U},\hat{N})$ $=\ln\Omega(\hat{U},\hat{N})$
$\tilde{\Psi}(\tfrac{1}{T},\hat{N})$ $=\ln\mathcal{Z}(\tfrac{1}{T},\hat{N})$
$\tilde{\Psi}(\tfrac{1}{T},\tfrac{\mu}{T})$ $=\ln\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T}),$
@@
$e^{\tilde{\Psi}(\hat{U},\hat{N})}$ $=e^{\tilde{S}(\hat{U},\hat{N})}=\Omega(\hat{U},\hat{N})$ (3)
$e^{\tilde{\Psi}(\tfrac{1}{T},\hat{N})}$ $=\sum_{\hat{U}}e^{\tilde{\Psi}(\hat{U},\hat{N})}e^{-\tfrac{1}{T}\hat{U}}$ (4)
$e^{\tilde{\Psi}(\tfrac{1}{T},\tfrac{\mu}{T})}$ $=\sum_{\hat{N}}\sum_{\hat{U}}e^{\tilde{\Psi}(\hat{U},\hat{N})}e^{-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}},$ (5)
where $\sum_{\hat{U}}$ and $\sum_{\hat{N}}$ represent the sum on all possible values of $\hat{U}$ and $\hat{N}$ (the sum is replaced by an integral for continuous variables). The mathematical transformation that generates partition functions (or SM Massieu functions $\tilde{\Psi}$) is the Laplace transform (see, for example, (Greiner et al., 1995, p. 247)). The fundamental approximation that connects thermodynamics and statistical mechanics is to assume that there is one term that is overwhelmingly larger than the others in the sums (or integrals) above; we have this term when $\hat{U}$ and $\hat{N}$ take their thermodynamic average values $U$ and $N$, so we can write
$e^{\Psi(U,N)}$ $=e^{S(U,N)}=\Omega(U,N)$ (6)
$e^{\Psi(\tfrac{1}{T},N)}$ $=e^{\Psi(U,N)}e^{-\tfrac{1}{T}U}$ (7)
$e^{\Psi(\tfrac{1}{T},\tfrac{\mu}{T})}$ $=e^{\Psi(U,N)}e^{-\tfrac{1}{T}U+\tfrac{\mu}{T}N},$ (8)
and we recover the definitions of TH Massieu functions generated by Legendre transforms (2). Since we have made an approximation, $\tilde{\Psi}$ and $\Psi$ are not exactly the same.
@@
$P(\omega)=\frac{\exp[-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}]}{\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})}=\exp[-\tilde{\Psi}(\tfrac{1}{T},\tfrac{\mu}{T})-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}].$ (9)
In order to obtain the probability of the macrostate $\hat{\vec{X}}=(\hat{U},\hat{N})$, we have to multiply by the corresponding number of compatible microstates $\Omega(\hat{U},\hat{N})=\exp[\tilde{\Psi}(\hat{U},\hat{N})]$; we have,
$P(\hat{U},\hat{N})=\exp[-\tilde{\Psi}(\tfrac{1}{T},\tfrac{\mu}{T})+\tilde{\Psi}(\hat{U},\hat{N})-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}].$ (10)
@@
$P(\hat{N})=\exp[-\tilde{\Psi}(\tfrac{1}{T},\tfrac{\mu}{T})+\tilde{\Psi}(\tfrac{1}{T},\hat{N})+\tfrac{\mu}{T}\hat{N}],$ (11)
@@
In general, for a given microstate $\omega$ the extensive variables take the values $\hat{\vec{X}}=(\hat{X}_{1},\dots,\hat{X}_{r})$; the list does not include extensive variables fixed by external constraints; variables without hat, $X_{i}$, are the thermodynamic averages. The probability of the microstate $\omega$ is given by (see, for example, (Callen, 1985, Ch. 19))
$P(\omega)=\exp[-\tilde{\Psi}(\vec{Y})-\vec{Y}\cdot\hat{\vec{X}}],$ (12)
where $\vec{Y}=(Y_{1},\dots,Y_{r})$ are entropic intensive parameters defined by $Y_{i}=\frac{\partial S(\vec{X})}{\partial X_{i}}$ and set by the reservoir. Multiplying by the number of microstates with $\hat{\vec{X}}$, given by $\Omega(\hat{\vec{X}})=\exp[\tilde{\Psi}(\hat{\vec{X}})]$, we have the probability of $\hat{\vec{X}}$:
$P(\hat{\vec{X}})=\exp[-\tilde{\Psi}(\vec{Y})+\tilde{\Psi}(\hat{\vec{X}})-\vec{Y}\cdot\hat{\vec{X}}].$ (13)
If we have $s$ extensive variables and $r-s$ intensive parameters, that we can write as $\vec{X}^{s}=(X_{1},\dots,X_{s})$ and $\vec{Y}^{r-s}=(Y_{s+1},\dots,Y_{r})$, then the SM Massieu function is the logarithm of the partition function in the corresponding ensemble:
$\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})=\ln\mathcal{Z}(\vec{X}^{s},\vec{Y}^{r-s}),$ (14)
where we have used the average values of the extensive variables (without hat), but the same expression holds for $\hat{\vec{X}}^{s}$ for a given microstate. Also, the SM Massieu function that depends on $s$ extensive variables and $r-s$ intensive parameters is given by $r-s$ Laplace transforms of $\tilde{\Psi}(\hat{\vec{X}})$:
$e^{\tilde{\Psi}(\hat{\vec{X}}^{s},\vec{Y}^{r-s})}$ $=$ (15)
$\sum_{\hat{X}_{s+1}}\cdots\sum_{\hat{X}_{r}}$ $e^{\tilde{\Psi}(\hat{\vec{X}})}e^{-Y_{s+1}\hat{X}_{s+1}-\dots-Y_{r}\hat{X}_{r}}.$
@@
$=e^{-\tilde{\Psi}(\vec{Y})+\tilde{\Psi}(\hat{\vec{X}}^{s},\vec{Y}^{r-s})-Y_{1}\hat{X}_{1}-\dots-Y_{s}\hat{X}_{s}},$ (16)
@@ III.1 $\hat{U},\hat{N},V$ system
Let us first consider a simple specific case: we want to calculate the difference $\tilde{\Psi}(\tfrac{1}{T},N)-\Psi(\tfrac{1}{T},N)$ for the system in contact to energy and particle reservoirs. That is, the difference between the statistical mechanics free energy, obtained from the partition function, and the thermodynamic free energy. It can be shown that both quantities are related through their derivatives. By definition, we have
$\frac{\partial\Psi(\tfrac{1}{T},N)}{\partial N}=-\frac{\mu}{T}.$ (17)
Even for small systems, it is usually assumed that the particle number $N$ is large enough so that thermodynamic potentials can be taken as continuous functions of $N$ (Hill, 1964, p. 6). On the other hand, we can calculate the average of the derivative of $\tilde{\Psi}(\tfrac{1}{T},\hat{N})$. Taking the logarithm of Eq. (11) we have,
$\tilde{\Psi}(\tfrac{1}{T},\hat{N})=\ln P(\hat{N})+\tilde{\Psi}(\tfrac{1}{T},\tfrac{\mu}{T})-\tfrac{\mu}{T}\hat{N}.$ (18)
@@
$\left\langle\frac{\partial\tilde{\Psi}(\tfrac{1}{T},\hat{N})}{\partial\hat{N}}\right\rangle$ $=\left\langle\frac{1}{P(\hat{N})}\frac{\partial P(\hat{N})}{\partial\hat{N}}-\frac{\mu}{T}\right\rangle$
@@
$\left\langle\frac{\partial\tilde{\Psi}(\tfrac{1}{T},\hat{N})}{\partial\hat{N}}\right\rangle=\frac{\partial\Psi(\tfrac{1}{T},N)}{\partial N}.$ (21)
The average can be written in terms of mean squared fluctuations if we expand $\frac{\partial\tilde{\Psi}(\tfrac{1}{T},\hat{N})}{\partial\hat{N}}$ in a Taylor series around the mean value $N$:
$\left\langle\frac{\partial\tilde{\Psi}(\tfrac{1}{T},\hat{N})}{\partial\hat{N}}\right\rangle=\frac{\partial\tilde{\Psi}(\tfrac{1}{T},N)}{\partial N}+\frac{1}{2}\frac{\partial^{3}\Psi(\tfrac{1}{T},N)}{\partial N^{3}}\langle(\Delta\hat{N})^{2}\rangle,$ (22)
where $\Delta\hat{N}=\hat{N}-N$, and terms $O(N^{-2})$ are neglected [the mean squared or higher order fluctuations are $O(N)$, and the $i$-th derivative is $O(N^{-i+1})$]; the third derivative of $\tilde{\Psi}$ was replaced by the third derivative of $\Psi$ (without tilde) since the difference contributes with a term $O(N^{-2})$. Combining the last two equations,
$\frac{\partial}{\partial N}[\tilde{\Psi}(\tfrac{1}{T},N)-\Psi(\tfrac{1}{T},N)]=-\frac{1}{2}\frac{\partial^{3}\Psi(\tfrac{1}{T},N)}{\partial N^{3}}\langle(\Delta\hat{N})^{2}\rangle.$ (23)
@@
$\langle(\Delta\hat{N})^{2}\rangle=-\frac{\partial N}{\partial(-\tfrac{\mu}{T})}=-\frac{1}{\frac{\partial(-\tfrac{\mu}{T})}{\partial N}}=-\frac{1}{\frac{\partial^{2}\Psi(\tfrac{1}{T},N)}{\partial N^{2}}},$ (24)
@@
$\frac{\partial^{3}\Psi(\tfrac{1}{T},N)}{\partial N^{3}}=\frac{1}{\langle(\Delta\hat{N})^{2}\rangle^{2}}\frac{\partial\langle(\Delta\hat{N})^{2}\rangle}{\partial N}.$ (25)
@@
$\frac{\partial\Delta\tilde{\Psi}(\tfrac{1}{T},N)}{\partial N}=-\frac{1}{2}\frac{\partial}{\partial N}\ln\langle(\Delta\hat{N})^{2}\rangle,$ (26)
where $\Delta\tilde{\Psi}(\tfrac{1}{T},N)=\tilde{\Psi}(\tfrac{1}{T},N)-\Psi(\tfrac{1}{T},N)$; and, integrating in $N$,
$\Delta\tilde{\Psi}(\tfrac{1}{T},N)=-\frac{1}{2}\ln\langle(\Delta\hat{N})^{2}\rangle+c,$ (27)
@@
If interactions among components of the system can be neglected, the TH or SM Massieu function takes its ideal value, $\Psi_{\text{id}}$ or $\tilde{\Psi}_{\text{id}}$. Assuming the ideal situation, Eq. (27) is
$\Delta\tilde{\Psi}_{\text{id}}(\tfrac{1}{T},N)=-\frac{1}{2}\ln\langle(\Delta\hat{N})^{2}\rangle_{\text{id}}+c,$ (28)
The excess Massieu function (TH or SM) is defined as the difference $\Psi_{\text{ex}}=\Psi-\Psi_{\text{id}}$ (with or without tilde). Doing subtraction between Eqs. (27) and (28), we obtain,
$\Delta\tilde{\Psi}_{\text{ex}}(\tfrac{1}{T},N)=-\frac{1}{2}\ln\frac{\langle(\Delta\hat{N})^{2}\rangle}{\langle(\Delta\hat{N})^{2}\rangle_{\text{id}}}.$ (29)

### v2

@@ I Introduction
We present a general approach to demonstrate that those differences can be written in terms of thermodynamic fluctuations, that, in turn, can be obtained from thermodynamic quantities. This means that the thermodynamic equation of state contains the information needed to calculate non-extensive terms of statistical mechanics quantities. Or, the other way around, non-extensive terms of statistical mechanics quantities have information about thermodynamic fluctuations. Our analysis is restricted to systems large enough so that finite size effects, analyzed in, for example, Hill (1964); Bedeaux et al. (2020), other than the logarithmic term, can be neglected.
The difference $O(\ln N)$ is present in the non-ideal and ideal (negligible interactions) cases. Excess extensive quantities are obtained when the ideal value is subtracted. It can be seen that thermodynamic and statistical mechanics excess quantities differ by a term that is $O(N^{0})$. For example, the excess free energy of a system of particles, calculated using the partition function, includes a non-extensive term $O(N^{0})$; and the excess chemical potential has a non-extensive term $O(N^{-1})$. This term was calculated using the Widom insertion formula in Ref. Siepmann et al. (1992), where the result is presented as a finite size correction to numerical calculation of the excess chemical potential in a system with periodic boundary conditions. An equivalent expression was derived with a different procedure in Ref. Di Muro and Hoyuelos (2021), where it was shown that the non-extensive term is relevant for the evaluation of transition rates and, in turn, of the diffusion coefficient. Actually, the relationship between self-diffusion and collective diffusion coefficients, known as one of the the Darken equations Darken (1948), was derived in Di Muro and Hoyuelos (2021) using the non-extensive term of the excess chemical potential. Here we present a general approach, valid for any Massieu function, that includes the mentioned results; that is, it is not restricted to the analysis of the free energy. It is generally more convenient to use Massieu functions instead of thermodynamic potentials in order to analyze fluctuations (see (Callen, 1985, Ch. 19)); there are simple relationships that connect them.
@@ II Basic equations and notation
where $\Omega(\vec{X})$ is the number of microstates that are consistent with the macrostate $\vec{X}$. Units such that $k_{B}=1$ are used in order to simplify the notation; in this context, temperature has energy units. We use a tilde in order to identify statistical mechanics quantities. The thermodynamic entropy, homogeneous of order 1 in $N$, is $S(\vec{X})$. The difference is $\tilde{S}(\vec{X})-S(\vec{X})=O(\ln N)$.
First, we present the definition of Massieu functions in their thermodynamic (TH) and statistical mechanics (SM) versions for a specific situation: a particle sub-system of volume $V$ that is part of a larger complex; the complement of the sub-system plays the role of a reservoir that imposes a temperature $T$ and a chemical potential $\mu$. Local fluctuations in $U$ and $N$ are present, and $V$ is a constant parameter by definition. We consider that the vector $\vec{X}=(U,N)$ contains the extensive variables, but there can be other parameters, like $V$ in this case, that are fixed by external constraints. TH Massieu functions are obtained from successive applications of Legendre transforms, starting from the entropy. For example:
$S(U,N)$
$S[\tfrac{1}{T},N]$ $=S(U,N)-\tfrac{1}{T}U$ (2)
$S[\tfrac{1}{T},\tfrac{\mu}{T}]$ $=S(U,N)-\tfrac{1}{T}U+\tfrac{\mu}{T}N.$
Different Massieu functions $S$ are identified by their variables between square brackets; the entropy, $S(U,N)$, with the variables in parenthesis, is the first Massieu function. We have that $S[\tfrac{1}{T},N]=-F/T$, where $F$ is the Helmholtz free energy, and $S[\tfrac{1}{T},\tfrac{\mu}{T}]=PV/T$, where $P$ is the pressure. A fundamental hypothesis of thermodynamics is that thermodynamic potentials, and TH Massieu functions, are homogeneous functions of order 1 in $N$; that is, they are extensive quantities.
Now we turn to a statistical mechanics description. We have to specify how to calculate Massieu functions in this context. Instead of the Legendre transform, the mathematical transformation that generates SM Massieu functions starting from the entropy is the Laplace transform. Quantities with tilde, in the statistical mechanics context, contain non-extensive terms; we are interested in, for example, the difference $\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-S[\tfrac{1}{T},\tfrac{\mu}{T}]$ at order $\ln N$.
The TH Massieu functions $S[\tfrac{1}{T},N]$ and $S[\tfrac{1}{T},\tfrac{\mu}{T}]$, equal to $-F/T$ and $PV/T$, correspond to the logarithm of the canonical and grand canonical partition functions, $\mathcal{Z}(\tfrac{1}{T},N)$ and $\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})$, respectively, as long as non-extensive terms are neglected in the thermodynamic limit. The SM Massieu functions are given by
$\tilde{S}(\hat{U},\hat{N})$ $=\ln\Omega(\hat{U},\hat{N})$
$\tilde{S}[\tfrac{1}{T},\hat{N}]$ $=\ln\mathcal{Z}(\tfrac{1}{T},\hat{N})$
$\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]$ $=\ln\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T}),$
@@
$e^{\tilde{S}(\hat{U},\hat{N})}$ $=\Omega(\hat{U},\hat{N})$ (3)
$e^{\tilde{S}[\frac{1}{T},\hat{N}]}$ $=\sum_{\hat{U}}e^{\tilde{S}(\hat{U},\hat{N})}e^{-\frac{1}{T}\hat{U}}$ (4)
$e^{\tilde{S}[\frac{1}{T},\frac{\mu}{T}]}$ $=\sum_{\hat{N}}\sum_{\hat{U}}e^{\tilde{S}(\hat{U},\hat{N})}e^{-\frac{1}{T}\hat{U}+\frac{\mu}{T}\hat{N}},$ (5)
where $\sum_{\hat{U}}$ and $\sum_{\hat{N}}$ represent the sum on all possible values of $\hat{U}$ and $\hat{N}$ (the sum is replaced by an integral for continuous variables). The mathematical transformation that generates partition functions (or SM Massieu functions) is the Laplace transform (see, for example, (Greiner et al., 1995, p. 247)). The fundamental approximation that connects thermodynamics and statistical mechanics is to assume that there is one term that is overwhelmingly larger than the others in the sums (or integrals) above; we have this term when $\hat{U}$ and $\hat{N}$ take their thermodynamic average values $U$ and $N$, so we can write
$e^{S(U,N)}$ $=\Omega(U,N)$ (6)
$e^{S[\frac{1}{T},N]}$ $=e^{S(U,N)}e^{-\frac{1}{T}U}$ (7)
$e^{S[\frac{1}{T},\frac{\mu}{T}]}$ $=e^{S(U,N)}e^{-\frac{1}{T}U+\frac{\mu}{T}N},$ (8)
and we recover the definitions of TH Massieu functions generated by Legendre transforms (2). Since we have made an approximation, $\tilde{S}$ and $S$ are not exactly the same.
@@
$P(\omega)=\frac{\exp[-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}]}{\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})}=\exp[-\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}].$ (9)
In order to obtain the probability of the macrostate $\hat{\vec{X}}=(\hat{U},\hat{N})$, we have to multiply by the corresponding number of compatible microstates $\Omega(\hat{U},\hat{N})=\exp[\tilde{S}(\hat{U},\hat{N})]$; we have,
$P(\hat{U},\hat{N})=\exp[-\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]+\tilde{S}(\hat{U},\hat{N})-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}].$ (10)
@@
$P(\hat{N})=\exp[-\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]+\tilde{S}[\tfrac{1}{T},\hat{N}]+\tfrac{\mu}{T}\hat{N}],$ (11)
@@
In general, for a given microstate $\omega$ the extensive variables take the values $\hat{\vec{X}}=(\hat{X}_{1},\dots,\hat{X}_{r})$; the list does not include extensive parameters fixed by external constraints; variables without hat, $X_{i}$, are the thermodynamic averages. The probability of the microstate $\omega$ is given by (see, for example, (Callen, 1985, Ch. 19))
$P(\omega)=\exp[-\tilde{S}[\vec{Y}]-\vec{Y}\cdot\hat{\vec{X}}],$ (12)
where $\vec{Y}=(Y_{1},\dots,Y_{r})$ are entropic intensive parameters defined by $Y_{i}=\frac{\partial S(\vec{X})}{\partial X_{i}}$ and set by the reservoir. Multiplying by the number of microstates with $\hat{\vec{X}}$, given by $\Omega(\hat{\vec{X}})=\exp[\tilde{S}(\hat{\vec{X}})]$, we have the probability of $\hat{\vec{X}}$:
$P(\hat{\vec{X}})=\exp[-\tilde{S}[\vec{Y}]+\tilde{S}(\hat{\vec{X}})-\vec{Y}\cdot\hat{\vec{X}}].$ (13)
If we have $s$ extensive variables and $r-s$ intensive parameters, that we can write as $\vec{X}^{s}=(X_{1},\cdots,X_{s})$ and $\vec{Y}^{r-s}=(Y_{s+1},\cdots,Y_{r})$, then the SM Massieu function is the logarithm of the partition function in the corresponding ensemble:
$\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]=\ln\mathcal{Z}(\vec{X}^{s},\vec{Y}^{r-s}),$ (14)
where we have used the average values of the extensive variables (without hat), but the same expression holds for $\hat{\vec{X}}^{s}$ for a given microstate. Also, the SM Massieu function that depends on $s$ extensive variables and $r-s$ intensive parameters is given by $r-s$ Laplace transforms of $\tilde{S}(\hat{\vec{X}})$:
$e^{\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]}$ $=$ (15)
$\sum_{\hat{X}_{s+1}}\cdots\sum_{\hat{X}_{r}}$ $e^{\tilde{S}(\hat{\vec{X}})}e^{-Y_{s+1}\hat{X}_{s+1}\dots-Y_{r}\hat{X}_{r}}.$
@@
$=e^{-\tilde{S}[\vec{Y}]+\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]-Y_{1}\hat{X}_{1}\dots-Y_{s}\hat{X}_{s}},$ (16)
@@ III.1 $\hat{U},\hat{N},V$ system
Let us first consider a simple specific case: we want to calculate the difference $\tilde{S}[\tfrac{1}{T},N]-S[\tfrac{1}{T},N]$ for the system in contact to energy and particle reservoirs. That is, the difference between the statistical mechanics free energy, obtained from the partition function, and the thermodynamic free energy, since $\tilde{S}[\tfrac{1}{T},N]=\ln\mathcal{Z}(\frac{1}{T},N)=-\tilde{F}/T$ and $S[\tfrac{1}{T},N]=-F/T$. It can be shown that both quantities are related through their derivatives. By definition, we have
$\frac{\partial S[\tfrac{1}{T},N]}{\partial N}=-\frac{\mu}{T}.$ (17)
Even for small systems, it is usually assumed that the particle number $N$ is large enough so that thermodynamic potentials can be taken as continuous functions of $N$ (Hill, 1964, p. 6). On the other hand, we can calculate the average of the derivative of $\tilde{S}[\tfrac{1}{T},\hat{N}]$. Taking the logarithm of Eq. (11) we have,
$\tilde{S}[\tfrac{1}{T},\hat{N}]=\ln P(\hat{N})+\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-\tfrac{\mu}{T}\hat{N}.$ (18)
@@
$\left\langle\frac{\partial\tilde{S}[\tfrac{1}{T},\hat{N}]}{\partial\hat{N}}\right\rangle$ $=\left\langle\frac{1}{P(\hat{N})}\frac{\partial P(\hat{N})}{\partial\hat{N}}-\frac{\mu}{T}\right\rangle$
@@
$\left\langle\frac{\partial\tilde{S}[\tfrac{1}{T},\hat{N}]}{\partial\hat{N}}\right\rangle=\frac{\partial S[\tfrac{1}{T},N]}{\partial N}.$ (21)
The average can be written in terms of mean squared fluctuations if we expand $\frac{\partial\tilde{S}[\frac{1}{T},\hat{N}]}{\partial\hat{N}}$ in a Taylor series around the mean value $N$:
$\left\langle\frac{\partial\tilde{S}[\tfrac{1}{T},\hat{N}]}{\partial\hat{N}}\right\rangle=\frac{\partial\tilde{S}[\tfrac{1}{T},N]}{\partial N}+\frac{1}{2}\frac{\partial^{3}S[\tfrac{1}{T},N]}{\partial N^{3}}\langle(\Delta\hat{N})^{2}\rangle,$ (22)
where $\Delta\hat{N}=\hat{N}-N$, and terms $O(N^{-2})$ are neglected [the mean squared or higher order fluctuations are $O(N)$, and the $i$-th derivative is $O(N^{-i+1})$]; the third derivative of $\tilde{S}$ was replaced by the third derivative of $S$ (without tilde) since the difference contributes with a term $O(N^{-2})$. Combining the last two equations,
$\frac{\partial}{\partial N}[\tilde{S}[\tfrac{1}{T},N]-S[\tfrac{1}{T},N]]=-\frac{1}{2}\frac{\partial^{3}S[\tfrac{1}{T},N]}{\partial N^{3}}\langle(\Delta\hat{N})^{2}\rangle.$ (23)
@@
$\langle(\Delta\hat{N})^{2}\rangle=-\frac{\partial N}{\partial(-\tfrac{\mu}{T})}=-\frac{1}{\frac{\partial(-\tfrac{\mu}{T})}{\partial N}}=-\frac{1}{\frac{\partial^{2}S[\frac{1}{T},N]}{\partial N^{2}}},$ (24)
@@
$\frac{\partial^{3}S[\tfrac{1}{T},N]}{\partial N^{3}}=\frac{1}{\langle(\Delta\hat{N})^{2}\rangle^{2}}\frac{\partial\langle(\Delta\hat{N})^{2}\rangle}{\partial N}.$ (25)
@@
$\frac{\partial\Delta\tilde{S}[\tfrac{1}{T},N]}{\partial N}=-\frac{1}{2}\frac{\partial}{\partial N}\ln\langle(\Delta\hat{N})^{2}\rangle,$ (26)
where $\Delta\tilde{S}[\tfrac{1}{T},N]=\tilde{S}[\tfrac{1}{T},N]-S[\tfrac{1}{T},N]$; and, integrating in $N$,
$\Delta\tilde{S}[\tfrac{1}{T},N]=-\frac{1}{2}\ln\langle(\Delta\hat{N})^{2}\rangle+c,$ (27)
@@
If interactions among components of the system can be neglected, the TH and SM Massieu functions take their ideal values, $S_{\text{id}}[\tfrac{1}{T},N]$ and $\tilde{S}_{\text{id}}[\tfrac{1}{T},N]$. Assuming the ideal situation, Eq. (27) is
$\Delta\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln\langle(\Delta\hat{N})^{2}\rangle_{\text{id}}+c,$ (28)
The excess Massieu function (TH or SM) is defined as the difference $S_{\text{ex}}=S-S_{\text{id}}$ (with or without tilde). Doing subtraction between Eqs. (27) and (28), we obtain,
$\Delta\tilde{S}_{\text{ex}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln\frac{\langle(\Delta\hat{N})^{2}\rangle}{\langle(\Delta\hat{N})^{2}\rangle_{\text{id}}}.$ (29)